# A Survey on 3D Gaussian Splatting Applications: Segmentation, Editing, and Generation

### v3

@@ 4.3 Performance Benchmarking: 3DGS Generation
$\bullet$ Metrics. Common evaluation metrics we used are listed below:
MLLM evaluation use multimodal large language models (e.g., GPT-4, LLaVA [306]) to evaluate generation quality from multiple dimensions. Common evaluation metrics include BLIP-VQA [307], CLIP-IQA [308], Q-Align [309] and others.
Natural Image Quality Evaluator (NIQE) computes image quality by extracting natural scene statistics from image patches, comparing them to a Gaussian model trained on pristine images, and calculating the deviation as a quality score.
Blind/Referenceless Image Spatial Quality Evaluator (BRISQUE) estimates perceptual quality by extracting natural scene statistics from locally normalized luminance and applying a regression model trained on human-rated distorted images.
Kernel Inception Distance measures distributional differences between real and generated images by computing maximum mean discrepancy with polynomial kernels on Inception-v3 features.
Chamfer distance computes the similarity between two point sets by computing the average closest-point distance in both directions.
Thresholded Symmetric Epipolar Distance (TSED) calculates the number of consistent frame pairs in a sequence.
Inception Score evaluates the quality of generated images using Inception-V3 [291].
CMMD [310] is based on richer CLIP embeddings and the maximum mean discrepancy with the Gaussian RBF kernel.
HPSv2 [311] is a scoring model that more accurately predicts human preferences for generated images.
User study. It evaluates results by gathering human feedback on preferences, appearance quality, and geometric quality.
@@
$\bullet$ Results. We select two representative datasets, GSO [247] and Objaverse [246], as benchmarks for 3DGS generation. The performance of various methods is summarized in Table VIII and Table IX, respectively. Due to the limited number of publicly comparable baselines, we provide a quantitative assessment based on the available results. On GSO dataset, NovelGS [235] achieves the best performance by tailoring the diffusion process to suit the 3D Gaussian representation. It supports both text-to-3D and image-to-3D generation, demonstrating strong generalization across input modalities. On Objaverse dataset, Atlas-Gaussians [225] performs best, leveraging latent space optimization to enhance generation quality. It primarily supports text-conditioned generation and shows promising results on open-domain content.
5 Future Directions
$\bullet$ Large-Scale Feed-Forward Learning. To overcome the inefficiency of per-scene optimization, especially under dense calibrated images conditions, more and more methods adopt feed-forward architectures to enable fast and generalizable segmentation, editing, and generation. In the future, future research should explore large-scale training regimes across diverse scenarios and domains. This includes scaling up data volume, improving data diversity (e.g., egocentric, indoor, outdoor), and designing robust pipelines that generalize beyond static, clean environments.
$\bullet$ Faithful Evaluation Metrics. Current evaluation protocols often inherit metrics from traditional 2D or NeRF-based tasks, which may not align well with 3DGS goals. Developing reliable and interpretable metrics specifically designed for evaluating application tasks from a 3D Gaussian perspective is essential to track progress and guide future research. Existing metrics mainly evaluate accuracy after projection onto 2D images, lacking meaningful insights into the inherent 3D geometric and semantic consistency.
$\bullet$ Synthesizing 3D Data. Given the persistent shortage of annotated 3D data, future research can leverage the generative capabilities of 3DGS to synthesize large-scale, high-quality 3D datasets from abundant 2D image collections. Such synthetic datasets could significantly enhance training data diversity and volume, ultimately driving substantial improvements in the performance of 3DGS application tasks. Additionally, combining synthesized data with domain adaptation techniques could further bridge the domain gap between synthetic and real-world data, enhancing the model’s robustness and generalization capabilities.
$\bullet$ Toward Generalist Models. Instead of designing separate models for segmentation, editing, and generation, a promising direction is to develop generalist 3DGS-based architectures that can perform multiple scene-level tasks within a unified framework. Developing such versatile models can facilitate the sharing of learned representations across tasks, resulting in improved efficiency, reduced redundancy in training, and enhanced generalization capabilities.
$\bullet$ Combining 3D Foundation Model. Emerging 3D foundation models such as VGGT [312] and FLARE [313] demonstrate robust, generalizable representations by pretraining on diverse 3D tasks, including depth estimation, point cloud reconstruction, camera prediction, and 3D point tracking from multi-view images. These models encode rich geometric priors that can benefit various downstream applications. Future work may incorporate them into 3DGS to improve generalization and efficiency.
6 Conclusion
This survey systematically explores recent advances in utilizing 3D Gaussian Splatting (3DGS) for downstream application tasks, representing a pioneering effort in summarizing this emerging field. We first introduce essential background knowledge, covering fundamental concepts and foundational models. Subsequently, we comprehensively review over 200 representative methods across several key tasks, including segmentation, editing, generation, and other related application tasks, providing a structured categorization from a technical perspective. Furthermore, we present a thorough analysis of existing benchmarks and conduct fair comparisons of representative models clearly. Lastly, we highlight current challenges and outline promising future directions to inspire continued research efforts toward enhanced high-level tasks.
Appendices
A.1 Dataset Statistics
Due to space limitations in the main manuscript, we present the full table of the statistics of representative 3DGS segmentation, editing, and generation datasets in Table A.1, Table A.2, and Table A.3.
TABLE A.1: Statistics of representative 3DGS segmentation datasets. See §4.1 for more detailed descriptions.
  Dataset Venue #Scene #Views(Avg.) Characterization
TABLE A.2: Statistics of representative 3DGS editing datasets. See §4.2 for more detailed descriptions.
  Dataset Venue #Scenes #Views(Avg.) Characterization
TABLE A.3: Statistics of representative 3DGS generation datasets. See §4.3 for more detailed descriptions.
  Dataset Venue #Type #Scenes Characterization
[…]

### v4

@@ A.1.3 Performance Benchmarking: 3DGS Generation
$\bullet$ Metrics. Evaluation in 3DGS generation usually involves several complementary aspects, including perceptual image quality, distributional realism, geometry-level quality, and human-centered evaluation.
Perceptual image quality. Natural Image Quality Evaluator (NIQE) [396] and Blind/Referenceless Image Spatial Quality Evaluator (BRISQUE) [397] are no-reference image quality metrics that estimate the naturalness of rendered results based on image statistics. They are useful for coarse perceptual assessment, but they do not explicitly measure semantic correctness or 3D consistency.
Distributional realism. Kernel Inception Distance (KID) [398], Inception Score [399], and CMMD [350] compare generated images against reference distributions in feature space. These metrics are more suitable for evaluating the realism and diversity of a set of generated images, but they may not directly reflect view consistency or geometric faithfulness.
Geometry-level quality. Chamfer Distance [400] measures similarity between generated and reference point sets, and is therefore more informative for geometric accuracy than purely image-based metrics. Thresholded Symmetric Epipolar Distance (TSED) [401] evaluates multi-view or sequential consistency by measuring the number of geometrically consistent frame pairs.
Semantic and human-centered evaluation. MLLM-based evaluation (e.g., GPT-4, LLaVA [352]) and HPSv2 [351] provide higher-level assessments of semantic quality or human preference, while user study remains important for judging appearance quality, geometric plausibility, and subjective preference in ways that automatic metrics may miss.
Overall, image-quality metrics, geometry metrics, and human-centered metrics capture different dimensions of generation performance. Strong scores on one class of metrics do not necessarily imply strong performance on the others, so a balanced evaluation typically requires several complementary criteria.
@@
$\bullet$ Results. We select two representative datasets, GSO [359] and Objaverse [358], as benchmarks for 3DGS generation. The performance of various methods is summarized in Table A.11 and Table A.12, respectively. Due to the limited number of publicly comparable baselines, we provide a quantitative assessment based on the available results. On GSO dataset, NovelGS [317] achieves the best performance by tailoring the diffusion process to suit the 3D Gaussian representation. It supports both text-to-3D and image-to-3D generation, demonstrating strong generalization across input modalities. On Objaverse dataset, Atlas-Gaussians [307] performs best, leveraging latent space optimization to enhance generation quality. It primarily supports text-conditioned generation and shows promising results on open-domain content.